# Learn, Check, Test – Security Testing Using Automata Learning and Model Checking

### v1

@@ Abstract
Cyber-physical systems are part of industrial systems and critical infrastructure. Therefore, they should be examined in a comprehensive manner to verify their correctness and security. At the same time, the complexity of such systems demands such examinations to be systematic and, if possible, automated for efficiency and accuracy. A method that can be useful in this context is model checking. However, this requires a model that faithfully represents the behavior of the examined system. Obtaining such a model is not trivial, as many of these systems can be examined only in black box settings due to, e.g., long supply chains or secrecy. We therefore utilize active black box learning techniques to infer behavioral models in the form of Mealy machines of such systems and translate them into a form that can be evaluated using a model checker. To this end, we will investigate a cyber-physical systems as a black box using its external communication interface. We first annotate the model with propositions by mapping context information from the respective protocol to the model using Context-based Proposition Maps (CPMs). We gain annotated Mealy machines that resemble Kripke structures. We then formally define a template, which we use to transfer the structures in to a format to be processed by a model checker. We further define generic security properties based on basic security requirements (authentication, confidentiality, privilege levels, and key validity). Due to the used CPMs, we can instantiate these properties with a meaningful context to check a specific protocol, which makes the approach flexible and scalable. Furthermore, the gained model can be easily altered to introduce non-deterministic behavior (like timeouts) or faults and examined if the properties still hold under these different conditions. Lastly, we demonstrate the versatility of the approach by providing case studies of very different systems (a passport and an automotive control unit), speaking different communication protocols (NFC and UDS), checked with the same tool chain and the same security properties.
@@ 1 Introduction
Cyber-physical systems in all domains of everyday life become more interconnected and complex. Various domains like automotive, production, but also critical infrastructures like electronic government systems naturally have to provide communication interfaces to provide more smart and usable functionality. It is therefore crucial to create methodologies to verify the correctness and security of these interfaces. Since those systems are mostly proprietary, verifying their security has become an increasingly laborious task that is afflicted with a high degree of uncertainty. Many approaches rely on static, pre-defined test suites that do not take any architectural or behavioral information into account. Therefore, validation of industrial systems and critical infrastructure often include the use of formal methods that provide more rigor in verification [1]. Such methods, including model checking, often come with the disadvantage of a high effort (especially with modeling systems) in conjunction with being hard to apply at black box systems. To mend these disadvantages, we try to automate the model creation using active automata learning. This technique allows for an automated model generation to be used in a checking process, and is also a technique that is destined for black box systems [2]. We therefore combine the static model checking with dynamic active learning. Furthermore, we want to be able to not only verify the systems, but also correct errors and simulate different behavior. We can also use this technique to create corner-cases for verification by altering the system behavior to a Maximum Credible Accident (MCA) scenario [3]. We therefore follow the approach to translate the learned models into a modeling language (see Sections 2.4 and 5.2).
@@ 1.1 Contribution
This paper combines formal methods, namely active automata learning of existing systems and model checking, to create a tool chain for checking cyber-physical systems (particularly communication protocols) for their correctness and security. Our main contributions are:
@@
    An approach for annotating learned models (Mealy machines [4]) with propositions using context-based proposition maps (CPMs - see Section 5.1), effectively creating Kripke structures [5] (combined into a single model with the Mealy machine as a hybrid annotated Mealy machine),
@@
    A formally defined and implemented template that allows for creating Rebeca models [6] from annotated Mealy machines,
@@
    Generalized properties to check security requirements along with a easily usable, scalable method for concretization using CPMs,
@@
    Two case studies of communication protocols (NFC [7] and UDS [8]) demonstrating the applicability and versatility of the approach.
@@ 1.2 Approach
We utilize active automata learning via the LearnLib Java library to infer Mealy machine models of a respective System-under-learning (SUL). We therefore created dedicated protocol adapters and defined input alphabets (Section 4). Since this model does not intrinsically contain any properties, we annotate it with sensible propositions. For that, we define using context-based proposition maps (CPMs) that define when propositions become true or false (Section 5.1). We created a tool that uses these maps to annotate Mealy machines represented in the GraphViz format, receiving an combined annotated Mealy machine. We then convert this structure into Rebeca code and verify it to be faithful to the original model (Sections 5.2 and 5.6). Once this step has been performed, we utilize the Rebeca Model Checking tool, RMC [9] to check the models for properties defined in linear temporal logic (LTL). These properties are partly specific to the scrutinized protocol, but can also partly be generalized to check very diverse protocols for the same properties. Eventually, if a property violation is found, the trace can be concretized via the learner’s interface and be used as a test case for verification on the live system. If this confirms the violation, that indicates an issue. Otherwise, the different behavior serves as a counterexample for the learner in a new iteration of the process. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. We automated the complete process by combining (a) the learner(s) using LearnLib [10] (Section 4), (b) a tool for generating Kripke structures and translating it into a Rebeca model to be checked (Section 5) and (c) the Rebeca model checking tool [9], used with generic security properties (Section 6).
The remainder of this paper outlines as follows: Section 2 contains basics, while Section 3 outlines general security requirements to be checked. Section 4 outlines the learner interfaces, Section 5 the Rebeca code generation including a formally defined template and a formal description of the CPM concept. Section 6 the model checking, including property definition. Section 7 briefly describes the relation to testing. Section 8 presents the case studies, Section 9 the related work and Section 10 concludes the paper.
@@ 2.1 Automata and Kripke Structures
Automata, or state machines, are a fundamental concept of computer science  [11, 12]. To describe cyber-physical systems, we use a certain type of automata, namely Mealy machines [4]. A Mealy machine $\mathcal{M}$ is defined as $\mathcal{M}=(Q,\Sigma,\Omega,\delta,\lambda,q_{0})$, with $Q$ being the set of states, $\Sigma$ the input alphabet, $\Omega$ the output alphabet (that may or may not be identical to the input alphabet), $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), and $q_{0}$ the initial state. A state machine can describe a system in a way that for any defined set of inputs (input word), it delivers the correct set of outputs (output word) [13]. While Mealy machines are widely used to model system behavior, Kripke structures [5] are basic structures used for model checking [14]. A Kripke structure is similar to a Mealy machine, as both can be seen as types of Labeled Transition Systems (LTS) [15, 16]. However, they differ in that instead of labeling the transitions, Kripke structures have their states labeled with propositions. Coming from propositional logic, these are logical attributes that can be true or false. Any state in which this attribute is true gains it as a proposition. Each state can have an element of the power set of all propositions. The formal definition of a Kripke structure ($\mathcal{K}$) is $\mathcal{K}=(S,S_{0},R,\mathcal{L})$ with $S$ being the set of states, $S_{0}\subseteq S$ the set of input states, $R\subseteq S\times S$ a transition relation and $\mathcal{L}=S\rightarrow 2^{AP}$ the labeling function that attributes an element of the power set of atomic propositions (AP{AP}) to the states.
@@ 2.2 Automata Learning
Automata learning was developed as a method to learn regular languages [17]. A learner (i.e., an algorithm) is allowed to ask two kinds of questions: membership queries and equivalence queries. We then assume a teacher that knows all of the rules that determine well-defined words of that language. Membership queries give examples of input words($\sigma\in\Sigma|W_{\sigma}=\langle\sigma_{1},\sigma_{2}..\sigma_{n}\rangle$), where the teacher answers wether the example is part of the language. After a sufficient amount of queries, the learner infers a hypothesis model of an automaton describing the language (in this case a Deterministic Finite Acceptor – DFA). It may then submit an equivalence query, where it presents the hypothesis to the teacher. The teacher either answers that the hypothesis is correct or presents a counterexample. To fit to real-world cyber-physical systems, approach has been adapted to learn Mealy machines as well, alongside with more efficient algorithms [18]. In this case, membership queries yield corresponding output words ($\omega\in\Omega|W_{\omega}=\langle\omega_{1},\omega_{2}..\omega_{n}\rangle$). One popular field to use automata learning is to learn and test black box systems. In this case there is no teacher available that knows the interna of the examined system (SUL). Therefore, membership queries are just passed on to the live system [^1] We are talking about active automata learning here. It is also possible to take the answers from recorded data in passive automata learning.. Equivalence queries are implemented by a sufficient amount of conformance tests [19]. If all are passed, the hypothesis is accepted of correct. Otherwise, a failed conformance test automatically poses a counterexample. In this work, we use Learnlib [10], a well proliferated Java library for automata learning. It provides a classes for interfacing with SULs, implementations of well-known algorithms (L*, Rivest-Schapire, AAAR, ADT, KV, DHC and TTT, as well as an addon for L#  [20]), and conformance testing strategies (complete depth-bounded exploration, random words, random walk, W-method, Wp-method). Other libaries are AALpy [21] and Libalf [22]).
@@ 2.2.1 Interface and Abstraction Layer
To interact with real-world systems, we not only need the input alphabet ($\Sigma$), but also a means to communicate with the SUL. To keep the as universal as possible, we therefore define an intermediate layer that serves two purposes: provide an interface to the SUL (including necessary error correction, timeout-handling, etc.) and abstracting the input layer. The latter, called mapper, is necessary to reduce the input space to a size feasible to be exhaustively explored. The mapper is formally defined as a state machine $A=(\Sigma,\Omega,Q,q_{0},\delta,X,Y,\nabla)$, with $\Sigma$ the abstract input alphabet, $\Omega$ the abstract output alphabet, $Q$ being the set of states, $q_{0}$ the initial state, $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), $X$ the concrete input alphabet, $Y$ the concrete output alphabet, and $\nabla$ the abstraction function ($\nabla:\Sigma\cup\Omega\rightarrow X\cup Y$. Inversely, the same mapper can convert the received concrete output from the SUL back to abstract output to present the learner by replacing the abstraction function with its inverse $\nabla^{-1}:X\cup Y\rightarrow\Sigma\cup\Omega$. The latter must not necessarily be fully defined in all cases. For many protocols, the learner is able to handle concrete outputs and to learn $\Omega$ on-the-fly. For these cases a partial mapper with $A=(\Sigma,Q,q_{0},\delta,X,\nabla_{in})$ is sufficient where $\nabla_{in}:\Sigma\rightarrow X$. However, for some concrete outputs ($Y\prime\subseteq Y$) the inverse mapper must still be defined ($\nabla_{out}^{-1}:Y\prime\rightarrow\Omega$), since some inputs provoke non-deterministic values in the sense that they would not yield the same output even with the exact equal input sequence, which a non-deterministic Mealy machine learner cannot handle. Examples of such outputs are timestamps, counters, and random numbers (such as nonces).
@@ 2.3 Linear Temporal Logic and Model Checking
Linear Temporal Logic (LTL) is an extension of propositional logic [23] that allows for expressing logical temporal [^2] In this context, temporal is not to be confused with timed. LTL only allows statements about the modality and succession order of events to occur, not about a duration of any kind. modalities [24]. That is, LTL combines atomic propositions (attributes of a state that can be true or false) formulas using logical operators and extends them with the following modalities:
@@ 2.4 Rebeca
Rebeca is a modeling language with model checking support specifically designed to model and model check cyber-physical (i.e., reactive) systems [6]. Rebeca resembles the Java programing language to ease its usage for engineers. It conceptually uses rebecs to model reactive actors. Rebecs are similar to Java objects, but possess interfaces (message servers) to interact with each other and contain a message queue to process messages from other Rebecs in a FIFO manner. Its IDE, Afra [9], contains a model checker, Modere [27]. Each Rebec can have state variables that can constitute checkable properties. The properties are defined as LTL formula in a separate property file. Apart from the usability for software engineering and modeling and model checking cyber-physical systems, it is possible to convert code into Lingua Franca [28]. Lingua Franca is a coordination language, providing a deterministic concurrency model, that supports C, C++, Python, TypeScript, and Rust as targets [29]. Eventually, a Rebeca code can produce a verified program in one of these languages via Lingua Franca. Formally, a Rebeca model $\mathcal{M}$ is defined as $\mathcal{M}=||_{i\in\mathcal{I}}\text{ }r_{i}$, which is the set of concurrent actors (also called rebec for reactive object) $r_{i}$, with $i\in\mathcal{I}$ being an identifier from the set of all identifiers in the model [30]. An actor is defined as $r_{i}:=\langle V_{i},M_{i},K_{i}\rangle$, where $i$ is a unique identifier, $Vi$ the set of state variables, $M_{i}$ the set of method identifiers (local methods and remotely callable methods called message servers), and $K_{i}$ the list of other actors known to actor $i$. Remote method (i.e., message server) calls between actors occur by sending messages ${msg}=\langle{sendid},i,{mtdid}\rangle$, with ${sendid}\in\mathcal{I}$ being the sender actor’s identifier, $i\in\mathcal{I}$ the receiver actor’s identifier, and $mtdid\in M_{i}$ the identifier of the method called on the sender side.

### v2

# arXiv:2509.22215v2[cs.CR] 27 Feb 2026

@@ Abstract
Security verification of communication protocols in industrial and safety-critical systems is challenging because implementations are often proprietary, accessible only as black boxes, and too complex for manual modeling. As a result, existing security testing approaches usually depend on incomplete test suites and/or require labor-intensive modeling, limiting coverage, scalability, and trust. This paper addresses the problem of systematically verifying protocol security-properties without access to internal system models. We propose a flexible and scalable method for formal verification of communication protocols that combines active automata learning with model checking to enable rigorous security analysis of black-box protocol implementations. Behavioral models are first inferred from system interactions using automata learning. We then propose Context-based Proposition Maps (CPMs) to enrich the learned models with semantic information, yielding annotated Mealy machines that bridge the gap between learned behavior and property verification. These machines are automatically transformed into verifiable models in the Rebeca modeling language, enabling model-checking of generic security properties such as authentication, confidentiality, privilege levels, and key validity, while allowing protocol-specific properties to be manually added. The CPMs are also used to populate these generic properties with definitive propositions. Furthermore, the gained model can be easily altered to introduce non-deterministic behavior (like timeouts or faults) and be examined if the properties still hold under these different conditions. As a result, we gain an automated tool chain that spans from model learning to security property checking of specific protocols. The key contributions include: (i) a method for augmenting learned protocol models with security-relevant propositions, (ii) a fully automated transformation pipeline from learned models to model-checking artifacts, (iii) reusable, generic security property templates that are instantiated in protocol-specific models, and (iv) empirical validation through case studies demonstrating applicability in different protocols and domains. The results show that the approach enables scalable and systematic discovery of security vulnerabilities in black-box systems while reducing modeling effort and improving automation compared with traditional verification workflows.
@@ 1 Introduction
Interconnected electronic systems in all domains of everyday life become more both ubiquitous and complex. Various domains like automotive, production, but also critical infrastructures like electronic government systems naturally have to provide communication capabilities to provide more smart and usable functionality. It is therefore crucial to create methodologies to verify the correctness and security of these systems. Often those systems are proprietary, which means that verifying their security has become an increasingly laborious task that is afflicted with a high degree of uncertainty. Many approaches rely on static, pre-defined test suites that do not take any architectural or behavioral information into account. Therefore, validation of industrial systems and critical infrastructure often include the use of formal methods that provide more rigor in verification [51]. Such methods, including model checking, often come with the disadvantage of a high effort (especially with modeling systems) in conjunction with being hard to apply at black box systems. To mend these disadvantages, we automate the model creation using active automata learning. This technique allows for an automated model generation to be used in a checking process, and is also a technique that is suitable for black box systems [55]. In this paper, we combine model checking with active learning, introducing novel concepts like Annotated Mealy Machines (AMMs) and Context-Based Proposition Maps (CPMs) to provide the missing link between those two components. Furthermore, our technique is able to not only verify the systems, but also help in correcting errors and simulate different behaviors. We can use the technique to create corner-cases for verification by altering the system behavior to a maximum credible accident scenario [35].
@@ 1.1 Contribution
This paper combines formal methods, namely active automata learning of existing systems and model checking, to create a tool chain for checking interactive systems (particularly their communication protocols) for correctness and security. In general, formally verifying protocols is a complex and labor-intense task. Both creating a faithful system model and sensible, meaningful security properties is time-consuming and error-prone. We significantly reduce this workload by automating the model generation through making automatically learned models usable for model checking and by providing templates for generic security properties that can easily be instantiated for concrete protocols. The core idea for both are context-based proposition maps (CPMs). CPMs are simple, tabular rule sets that are used to define when certain propositions become true or false by actions on the transitions from one state to another. For the modeling, we use CPMs to annotate Mealy machines [31], obtained via automata learning, with checkable propositions. The result are hybrid structures between Mealy machines and Kripke structures we call annotated Mealy machines (AMMs). Creating a formally defined template, also provide an automated translation of AMMs into the Rebeca modeling language, where we can run it through a model checker. For the property definition, we provide generic properties (particularly authentication, confidentiality, privilege levels, and key validity) which are represented by composition of propositions that we also define through CPMs. If the protocol is known, CPMs are easy to create. The only task is to denote which actions in a protocol changes a proposition to true or false. This action is represented by an input or output as a label on a transition. We call this change of value of a proposition as gain or loss of the proposition in a state when the propositions becomes true or false, respectively. In summary, our main contributions are:
@@
    Proposing context-based proposition maps (CPMs) to be used for annotating the generic models and populating the generic properties with protocol-specific propositions,
@@
    Offering an approach for automatically annotating learned models (Mealy machines) with propositions using CPMs and creating annotated Mealy machines (AMMS) – hybrid structures of Mealy machines and Kripke structures (suitable for model checking),
@@
    Defining and implementing a template that allows for fully automatically creating Rebeca models from AMMs,
@@
    Defining generalized properties to check security requirements along with an easily usable and scalable method for deriving concrete properties based on the generalized ones using CPMs,
  • 5.
    Demonstrating the applicability and versatility of the approach using two case studies of communication protocols (NFC [17] and UDS [19]).
@@ 1.2 Approach
We utilize active automata learning via the LearnLib Java library to infer Mealy machine models of a respective System-under-learning (SUL). We therefore created dedicated protocol adapters and defined input alphabets (Section 3). Since this model does not intrinsically contain any properties, we annotate it with sensible propositions, creating a hybrid automaton of a Mealy machine and a Kripke structure, which we call an annotated Mealy machine (AMM). For that, we define context-based proposition maps (CPMs) that determine the ruleset when propositions become true or false (Section 4). We created a tool that uses these maps to annotate Mealy machines represented in the GraphViz format, receiving an AMM. We then convert this structure into Rebeca code and verify it to be faithful to the original model in an automated way, formally defining a Rebeca template for that purpose (Sections 6 and 6.5) . Once this step has been performed, we utilize the Rebeca Model Checking tool, RMC [23] to check the models for properties defined in linear temporal logic (LTL). These properties are partly specific to the scrutinized protocol, but can also partly be generalized to check very diverse protocols for the same properties: we define commonly important properties like authentication, confidentiality, privilege levels, and key validity (Section 7.2), while each protocol might still have constraints that are unique to it (like the order of security-related commands). Eventually, if a property violation is found, the trace can be concretized via the learner’s interface and be used as a test case for verification on the live system. If this confirms the violation, that indicates an issue. Otherwise, the different behavior serves as a counterexample for the learner in a new iteration of the process. That way, we implemented all the phases for going through learning, checking, and testing real-world examples. We automated the complete process by combining (a) the learner(s) using LearnLib [21] (Section 3), (b) a tool that generates an annotated Mealy machine and translating it into a Rebeca model to be checked (Section 6) and (c) the Rebeca model checking tool [23], used with generic security properties. Figure 1 provides an overview of the proposed automated process of model model checking learned Mealy machines for generic security properties.
Figure 1: Overview of the proposed approach; boxes denote artifacts and ellipses denote processes or tools. The items in amber are artifacts and processes defined in this paper, while boxes in blue are artifacts automatically generated by algorithms defined in this paper. Cyan are other items, namely: Mealy machines of SuTs, which are obtained via automata learning and the model checking result which is obtained by running RMC with the respective inputs. RMC is shown as an ellipse in Cyan. Generic security properties and Rebeca templates are one-time work. For each specific problem, only CPMs have to be defined individually.
The remainder of this paper outlines as follows: Section 2 contains basics. Section 3 outlines the automata learner details, Section 6 the Rebeca code generation including a formally defined template and a formal description of the CPM concept. Section 7 the model checking, including property definition. Section 8 briefly describes the relation to testing. Section 9 presents the case studies, Section 10 the related work and Section 11 concludes the paper.
@@ 2.1 Automata and Kripke Structures
Automata, or state machines, are a fundamental concept of computer science  [52, 57]. To describe reactive systems, we use a certain type of automata, namely Mealy machines [31]. A Mealy machine $\mathcal{M}$ is defined as $\mathcal{M}=(Q,\Sigma,\Omega,\delta,\lambda,q_{0})$, with $Q$ being the set of states, $\Sigma$ the input alphabet, $\Omega$ the output alphabet (that may or may not be identical to the input alphabet), $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), and $q_{0}$ the initial state. A state machine can describe a system in a way that for any defined set of inputs (input word), it delivers the correct set of outputs (output word) [39]. While Mealy machines are widely used to model system behavior, Kripke structures [24] are basic structures used for model checking [8]. A Kripke structure is similar to a Mealy machine, as both can be seen as types of Labeled Transition Systems (LTS) [50, 7]. However, they differ in that instead of labeling the transitions, Kripke structures have their states labeled with propositions. Coming from propositional logic, these are logical attributes that can be true or false. Any state in which this attribute is true gains it as a proposition. Each state can have an element of the power set of all propositions. The formal definition of a Kripke structure ($\mathcal{K}$) is $\mathcal{K}=(S,S_{0},R,\mathcal{L})$ with $S$ being the set of states, $S_{0}\subseteq S$ the set of input states, $R\subseteq S\times S$ a transition relation and $\mathcal{L}=S\rightarrow 2^{P}$ the labeling function that attributes an element of the power set of (atomic) propositions ($p\in P$}) to the states.
@@ 2.2 Automata Learning
Automata learning, more precisely active automata learning, was developed as a method to learn regular languages [2]. A learner (i.e., an algorithm) is allowed to ask two kinds of questions: membership queries and equivalence queries. We then assume a teacher that knows all of the rules that determine well-defined words of that language. Membership queries give examples of input words($\sigma\in\Sigma|W_{\sigma}=\langle\sigma_{1},\sigma_{2}..\sigma_{n}\rangle$), where the teacher answers whether the example is part of the language. After a sufficient amount of queries, the learner infers a hypothesis model of an automaton describing the language (in this case a Deterministic Finite Acceptor – DFA). It may then submit an equivalence query, where it presents the hypothesis to the teacher. The teacher either answers that the hypothesis is correct or presents a counterexample. To fit to real-world systems, approach has been adapted to learn Mealy machines as well, alongside with more efficient algorithms [54]. In this case, membership queries yield corresponding output words ($\omega\in\Omega|W_{\omega}=\langle\omega_{1},\omega_{2}..\omega_{n}\rangle$). One popular field to use automata learning is to learn and test black box systems. In this case there is no teacher available that knows the interna of the examined system (SUL). Therefore, membership queries are just passed on to the live system. Equivalence queries are implemented by a sufficient amount of conformance tests [37]. If all are passed, the hypothesis is accepted of correct. Otherwise, a failed conformance test automatically poses a counterexample. In this work, we use Learnlib [21], a well proliferated Java library for automata learning. It provides classes for interfacing with SULs, implementations of well-known algorithms (L*, Rivest-Schapire, AAAR, ADT, KV, DHC and TTT, as well as an addon for L#  [25]), and conformance testing strategies (complete depth-bounded exploration, random words, random walk, W-method, Wp-method). Other libaries are AALpy [33] and Libalf [6]).
@@ 2.2.1 Interface and Abstraction Layer
To interact with real-world systems, we not only need the input alphabet ($\Sigma$), but also a means to communicate with the SUL. To keep it as universal as possible, we therefore define an intermediate layer that serves two purposes: provide an interface to the SUL (including necessary error correction, timeout-handling, etc.) and abstracting the input layer. The latter, called mapper, is necessary to reduce the input space to a size feasible to be exhaustively explored. The mapper is formally defined as a state machine $A=(\Sigma,\Omega,Q,q_{0},\delta,X,Y,\nabla)$, with $\Sigma$ the abstract input alphabet, $\Omega$ the abstract output alphabet, $Q$ being the set of states, $q_{0}$ the initial state, $\delta$ the transition function ($\delta:Q\times\Sigma\rightarrow Q$), $\lambda$ the output function ($\lambda:Q\times\Sigma\rightarrow\Omega$), $X$ the concrete input alphabet, $Y$ the concrete output alphabet, and $\nabla$ the abstraction function ($\nabla:\Sigma\cup\Omega\rightarrow X\cup Y$. Inversely, the same mapper can convert the received concrete output from the SUL back to abstract output to present the learner by replacing the abstraction function with its inverse $\nabla^{-1}:X\cup Y\rightarrow\Sigma\cup\Omega$. The latter must not necessarily be fully defined in all cases. For many protocols, the learner is able to handle concrete outputs and to learn $\Omega$ on-the-fly. For these cases a partial mapper with $A=(\Sigma,Q,q_{0},\delta,X,\nabla_{in})$ is sufficient where $\nabla_{in}:\Sigma\rightarrow X$. However, for some concrete outputs ($Y\prime\subseteq Y$) the inverse mapper must still be defined ($\nabla_{out}^{-1}:Y\prime\rightarrow\Omega$), since some inputs provoke non-deterministic values in the sense that they would not yield the same output even with the exact equal input sequence, which a non-deterministic Mealy machine learner cannot handle. Examples of such outputs are timestamps, counters, and random numbers (such as nonces).
@@ 2.3 Linear Temporal Logic and Model Checking
Linear Temporal Logic (LTL) is an extension of propositional logic [41] that allows for expressing logical temporal modalities [4]. In this context, temporal is not to be confused with timed. LTL only allows statements about the modality and succession order of events to occur, not about a duration of any kind. That is, LTL combines atomic propositions (attributes of a state that can be true or false) formulas using logical operators and extends them with the following modalities:
@@ 2.4 Rebeca
Rebeca is a modeling language with model checking support specifically designed to model and model check reactive (i.e., communicating) systems [49]. Rebeca resembles the Java programing language to ease its usage for engineers. It conceptually uses rebecs to model reactive actors. Rebecs are similar to Java objects,
but possess interfaces (message servers) to interact with each other and contain a message queue to process messages from other Rebecs in a FIFO manner. Its IDE, Afra [23], contains a model checker, Modere [22]. Each Rebec can have state variables that can constitute checkable properties. The properties are defined as LTL formula in a separate property file. Apart from the usability for software engineering and modeling and model checking reactive systems, it is possible to convert code into Lingua Franca [46]. Lingua Franca is a coordination language, providing a deterministic concurrency model, that supports C, C++, Python, TypeScript, and Rust as targets [32]. Eventually, a Rebeca code can produce a verified program in one of these languages via Lingua Franca. Formally, a Rebeca model $\mathcal{R}$ is defined as $\mathcal{R}=||_{i\in\mathcal{I}}\text{ }r_{i}$, which is the set of concurrent actors (also called rebec for reactive object) $r_{i}$, with $i\in\mathcal{I}$ being an identifier from the set of all identifiers in the model [48]. An actor is defined as $r_{i}:=\langle V_{i},M_{i},K_{i}\rangle$, where $i$ is a unique identifier, $Vi$ the set of state variables, $M_{i}$ the set of method identifiers (local methods and remotely callable methods called message servers), and $K_{i}$ the list of other actors known to actor $i$. Remote method (i.e., message server) calls between actors occur by sending messages ${msg}=\langle{sendid},i,{mtdid}\rangle$, with ${sendid}\in\mathcal{I}$ being the sender actor’s identifier, $i\in\mathcal{I}$ the receiver actor’s identifier, and $mtdid\in M_{i}$ the identifier of the method called on the sender side.